# Rydberg Atomic Quantum Receivers for Classical Wireless Communication and Sensing

### v1

@@ Abstract
The Rydberg atomic quantum receiver (RAQR) is an emerging quantum precision sensing platform designed for receiving radio frequency (RF) signals. It relies on creation of Rydberg atoms from normal atoms by exciting one or more electrons to a very high energy level, which in turn makes the atom sensitive to RF signals. The RAQR realizes RF-to-optical conversion based on light-atom interaction relying on the so called electromagnetically induced transparency (EIT) and Aulter-Townes splitting (ATS), so that the desired RF signal can be read out optically. The large dipole moments of Rydberg atoms associated with rich choices of Rydberg states and various modulation schemes facilitate an ultra-high sensitivity ($\sim$ nV/cm/$\sqrt{\text{Hz}}$) and an ultra-broadband tunability (near direct-current to Terahertz). RAQRs also exhibit compelling scalability and lend themselves to the construction of innovative, compact receivers. Initial experimental studies have demonstrated their capabilities in classical wireless communications and sensing. To fully harness their potential in a wide variety of applications, we commence by outlining the underlying fundamentals of Rydberg atoms, followed by the principles, structures, and theories of RAQRs. Finally, we conceive Rydberg atomic quantum single-input single-output (RAQ-SISO) and multiple-input multiple-output (RAQ-MIMO) schemes for facilitating the integration of RAQRs with classical wireless systems, and conclude with a set of potent research directions.
@@ I Introduction
Next-generation (NG) wireless systems are expected to support a variety of functionalities and a multitude of applications. Integrated communication and sensing functionalities of NG wireless require particularly high sensitivity and data rates, while supporting large-scale connectivity and ultra-low latency. Therefore, the receivers should possess broadband tunability and wideband processing capability. Based upon advanced antennas and integrated circuit technologies, cutting-edge radio frequency (RF) receivers support RF signal reception by well-calibrated antennas, filters, amplifiers, and mixers [1]. The sensitivity is limited by the combination of the receiver’s bandwidth, noise figure, and the minimum signal-to-noise ratio (SNR) required to demodulate the desired signal. Given a specific sensitivity, strong RF reception capability can be attained by using multi-antenna based multiple-input multiple-output (MIMO) schemes and sophisticated signal processing methods to combat channel fading and noise contamination. To receive wideband RF signals at different carrier frequencies, both broadband tunability and wideband processing capability may be realized by stacking several integrated circuit branches along with extra metallic components, where each branch is responsible for a specific frequency band. Such RF systems are however band-limited and bulky, power-thirsty, and suffer from numerous design challenges, especially for large-scale, high-frequency receivers.
@@
The Rydberg atomic quantum receiver (RAQR) concept emerges as a radical solution to these RF reception challenges [2, 3]. It relies on the creation of Rydberg states (excitation), which have remarkably high dipole moment and hence are eminently suitable for sensing the impinging RF signals. The response to an RF signal in RAQR is detected optically via change in the atomic spectroscopy signal. The RF field’s amplitude measurements are directly linked to the International System of Units (SI) and they set absolute (atomic) standards for the RF electric fields from a metrology standpoint. Additionally, they are capable of directly down-converting RF signals, spanning from kilohertz (kHz) to Terahertz (THz), to the baseband without using any mixers, hence significantly simplifying the receiver’s structure and yielding a compact form-factor relative to conventional RF receivers.
@@
To fully unlock the potential of RAQRs in communications and sensing, we offer a panoramic overview of RAQRs. We commence by highlighting the fundamentals of Rydberg atoms and proceed to powerful receivers, including their principles, structures, theories, impairments, and their critical comparisons to RF receivers. We then summarize their capabilities and distinctive features, surveying typical studies from diverse research communities. Furthermore, we conceive an Rydberg atomic quantum single-input single-output (RAQ-SISO) system and a potential RAQ-MIMO scheme to facilitate the integration of RAQRs with existing classical wireless systems. We finally forecast several promising future directions in this very first article on RAQRs in wireless communication and sensing from an engineering perspective.
@@
Figure 1: (a) Energy level diagram showing sub-levels of an atom. (b) Scaling of the relevant properties of Rydberg states with the alkali atom principal number ($n$). (c) Radial probability distribution of Rydberg states of Cs atoms for $n=10,90$. (d) Electron probability distribution $|\psi_{nlm}|^{2}$ with $n=70$, $l=50$, $m=40$, for hydrogen-like atoms.
@@ II Rydberg Atomic Quantum Receiver
RARQs exploit the properties of Rydberg atoms, hence we commence by their brief overview, followed by their principles, structures and theories. Finally, we present the impairments of RAQRs, and compare them to classical RF receivers.
@@ II-A1 Atomic Structure [4]
Quantum mechanics provides a comprehensive description of the structure, properties and behavior of atoms. Each atom has a nucleus at its core and there are electrons that orbit around the nucleus. The energy of electrons in these orbits is quantized, i.e., they occupy only certain discrete energy-levels and cannot be traced in between these levels. However, electrons can transition from one energy level to another by absorbing or emitting a photon with energy equal to the difference between the energy levels. Electrons have a wave-particle dual nature. In the quantum mechanical description, the state of an electron is described by a wavefunction $\psi$, which contains all information about the electron’s position, energy, spin, orientation, etc. The square of the wavefunction, $|\psi|^{2}$ gives the probability density (i.e. probability per unit volume) of finding the electron in a certain region of space around the nucleus. These regions are termed as orbitals. The unique quantum state of an electron is described by four quantum numbers:
@@
  • Azimuthal or orbital angular momentum quantum number $\bm{l}$ (non-negative integer $l\leq n-1$): It accounts for the sub-shell (sub-level) that the electron occupies, with each sub-shell having unique characteristic shapes, as described in [4]. The letters $S,P,D,F,G,H,I,...$ are assigned to $l=0,1,2,3,4,5,6,...$ respectively.
@@
The energy level structure with orbitals up to $6P$ is shown in Fig. 1(a). If mediated by photons of electromagnetic (EM) field, the transition of an electron from one quantum state to another is called an electric dipole transition. Based on the conservation of angular momentum and parity [^1] Parity refers to the symmetry of a wavefunction under spatial inversion., only those (electric dipole) transitions are allowed which satisfy the selection rules: ① $\Delta l=\pm 1$, ② $\Delta m=0,\pm 1$, and ③ the initial and final states must have opposite parities.
@@ II-A2 Properties of Rydberg State
Alkali atoms have single valence electrons in their ground state in the outermost orbital. This represents a relatively loose bound and therefore it can be excited to a high energy level associated with a high principal quantum number, of say $n>10$ for atomic systems. In such a situation, the atom is a Rydberg atom and the corresponding atomic state is a Rydberg state. The Rydberg states of any alkali atom and a hydrogen atom have essentially similar structures and properties since in both cases there is a single electron occupying a high-$n$ state and having an effective core of unit positive charge at the center. The only difference lies in the finite size of the positively charged core, which is larger for alkali atoms than for hydrogen atoms. This leads to a small correction (via quantum the defect parameter $\delta_{l}$) to the energies for non-hydrogenic atoms in particular for low $l$ states. But again, one can say that the general properties of all Rydberg atoms are quite similar. The scaling of the relevant properties of Rydberg states with $n$ is shown in Fig. 1(b). The $n$-dependence arises from the Rydberg atom wavefunction characteristics. For RAQRs, the typical choice of alkali atoms falls on either Caesium (Cs) or Rubidium (Rb) due to the availability of commercial laser systems for the transition wavelengths of these species.
The Rydberg atoms are large, since the average distance from the nucleus $\langle r\rangle$ scales with $n^{2}$, resulting in large dipole matrix elements. This point is illustrated in Fig. 1(c), where the radial probability distribution $|rR_{nl}(r)|^{2}$ for state $nS$ of Cs atoms is plotted against $r$ (distance from nucleus) for increasing $n$. The atomic size increases dramatically with $n$ and its value approaches 1 $\mu m$ for $n=100$, while the ground-state atom is just 0.1 nm in size. The spatial extent and particular shape of the electron probability distribution $|\psi_{nlm}|^{2}$, for $n=70$, $l=50$, $m=40$, is captured in Fig. 1(d) for hydrogen atom. The Rydberg-atom wavefunction (with same quantum numbers) for any alkali atom would closely resemble to this electronic distribution for such high $l$ and $m$ states.
@@ II-A3 Rydberg Atoms in Electric Fields
The sensitivity of Rydberg atoms to electric fields arises due to the high polarizability of Rydberg states. In the presence of electric field, the energy levels shift for different states in a way similar to the splitting of hydrogen atoms. This is called the Stark effect and the associated Stark shifts can be calculated by direct diagonalization of the Hamiltonian matrix. Fig. 2 shows the Stark map of Cs atoms in the vicinity of the $30D$ state. The red dashed line shows the energy level in the absence of electric field. The low-$l$ ($l=0,1,2$) Rydberg states are non-degenerate and separated from the associated hydrogenic manifold. At low electric fields $\mathcal{E}$, the Stark shift is a second-order energy shift and scales quadratically as $\Delta E_{\text{DC}}=\alpha\mathcal{E}^{2}/2$, where $\alpha$ is the DC polarizability of the Rydberg state. For $l>3$, the states are degenerate and therefore experience a linear Stark shift similar to the hydrogen Stark spectrum.
When the electric field interacting with atoms is constant or fluctuates at a low frequency, it is called DC Stark shift. When a high-frequency electric field (RF field) interacts with the atoms, it results in AC Stark shift, which has the same quadratic scaling with the field amplitude $\mathcal{E}$. The magnitude and sign of the AC Stark shift depends on the detuning $\Delta$ of the RF field from the resonant transition frequency and scales as $\Delta E_{\text{AC}}\propto 1/\Delta$. This implies that the magnitude of the shift is larger when the RF electric field is closer to resonance, while its sign is positive (negative) for blue-detuned (red-detuned) RF fields. The Stark effect causes coupling between closely-spaced energy levels leading to the mixing of energy eigenstates.
II-B Principles, Structures and Theories of RAQRs
@@ II-B1 Rydberg-Electromagnetically Induced Transparency (EIT) Spectroscopy and the Standard Scheme[3]
EIT is a quantum interference phenomenon, in which two excitation pathways of a three-level system interfere to open a transmission window in the absorption spectrum of atomic resonance. This is typically accomplished by counter-propagating a pair of laser beams, each of which corresponds to an excitation pathway. Since this is a coherent process, it is ultra-sensitive to phase perturbations and energy level shifts of the three-level system.
@@
Rydberg excitation is prepared via a two-photon excitation process, where a pair of counter-propagating laser beams of different wavelengths, namely the ‘probe’ and ‘coupling (control)’, are overlapped inside an alkali vapor cell, as shown in Fig. 3(a). This is accomplished via EIT in a ladder configuration (see Fig. 3(b)). The probe beam excites the atom from its ground state to an intermediate state, while the coupling beam takes it to a high-energy (principle number $n$) state, namely Rydberg state. When an external RF signal’s frequency is (near) resonant with two adjacent Rydberg states, changes can be observed in the optical spectroscopic signal. As the RF field strength increases from low to high, the Rydberg-EIT transparency peak becomes lower (due to small absorption) and then spilts into two peaks. This is the Autler-Townes splitting (ATS) phenomenon. The corresponding changes observed in the probe signal induced by the addition of a coupling laser and RF field, have been illustrated in Fig. 3(c).
@@
This is the standard scheme (structure) used in RAQR. The electron transitions take place among the four energy levels (quantum states) |$\ket{1}$, |$\ket{2}$, |$\ket{3}$, and |$\ket{4}$ with resonant transition frequencies $\omega_{(P,C,s)}$ and detunings $\Delta_{(P,C)}$, as shown in Fig. 3(b). Each quantum state has its own decay rate, denoted by $\gamma$. The transition |$\ket{1}\rightarrow\ket{2}$ is typically chosen to be D2 transition of Cs (Rb). |$\ket{3}$ could be any Rydberg state, and the excitation is created by tuning the coupling laser wavelength to the appropriate value. Finally, |$\ket{4}$ is the neighboring Rydberg state and its choice is governed by the particular RF frequency that the RAQR would be set up to receive. In this model, the measurement sensitivity approaches the photon shot noise limit, which is $\sim$ µV/cm/$\sqrt{\text{Hz}}$ [3].
II-B2 Superheterodyne Scheme for Enhanced Sensitivity [5, 6, 7, 8]
RAQR’s measurement sensitivity can be further improved by exploiting the superheterodyne principle, where an auxiliary RF signal, namely the local oscillator (LO), is added to the structure to detect relatively weak RF signals. This scheme is presented in Fig. 3(d). The RF signal is typically dozens or hundreds of kHz away from the strong LO in this configuration. The LO facilitates the creating of microwave-dressed Rydberg states and in the presence of two RF signals the frequency splitting measurement is converted into a modulated amplitude measurement, since the Rydberg atoms act as an RF mixer. The LO and the desired RF signal having frequencies $\omega_{(L,s)}$ and detunings of $\Delta_{(L)}$ as well as $\delta_{s}$ (with respect to LO) respectively are shown in Fig. 3(e). The sensitivity has been shown to be in the range of 12.5-55 nV/cm/$\sqrt{\text{Hz}}$ [6]. This method improves the sensitivity and additionally provides access to both amplitude and phase sensing. The phase of the RF signal can be extracted from the output modulated signal, since the phase of LO is known. Similarly, the angle of arrival (AoA) of an RF signal [7] can be accurately determined. When this scheme is extended to multilevel configuration, it allows weak RF signal detection over a wide frequency range of $>$ 1 GHz [8].
In both of the schemes above, a balanced photodetector (at the position of ‘Photodetector’ in Fig. 3) is commonly used to eliminate the laser intensity noise and any noise coming from the atoms in the cell. In the superheterodyne structure, the signal gleaned from balanced photodetection is sent to the lock-in amplifier of Fig. 3(d) for demodulation. Further signal processing may then be performed for accurately inferring the RF amplitude, frequency, phase, and so on.
Figure 3: (a) Standard model of RAQRs and its (b) energy level diagram. (c) Probe laser transmission as a function of probe detuning for three cases - no coupling laser, coupling laser present, both coupling laser and RF field present. (d) Superheterodyne model of RAQRs and its (e) the energy level diagram. High dipole moment of Rydberg atoms is symbolically highlighted in (a) and (d).
II-B3 Theories of RAQRs
The four-level atomic system and its response to electric field can be theoretically simulated by solving Lindblad master equation [6]. The steady state solutions provide the long-term behavior of the system when it is interacting with the environment. The macroscopic quantities of interest can be obtained from these solutions. The different properties of Rydberg atoms may then be studied with the help of the Alkali Rydberg Calculator package[9].

### v2

@@ Abstract
Rydberg atomic quantum receivers (RAQRs) are emerging quantum precision sensing platforms designed for receiving radio frequency (RF) signals. It relies on creation of Rydberg atoms from normal atoms by exciting one or more electrons to a very high energy level, thereby making the atom sensitive to RF signals. RAQRs realize RF-to-optical conversions based on atom-light interactions relying on the so called electromagnetically induced transparency (EIT) and Autler–Townes splitting (ATS), so that the desired RF signal can be read out optically. The large dipole moments of Rydberg atoms associated with rich choices of Rydberg states facilitate an ultra-high sensitivity ($\sim$ nV/cm/$\sqrt{\text{Hz}}$) and an ultra-broadband tunability (direct-current to Terahertz). RAQRs also exhibit compelling scalability and lend themselves to the construction of innovative, compact receivers. Initial experimental studies have demonstrated their capabilities in classical wireless communications and sensing. To fully harness their potential in a wide variety of applications, we commence by outlining the underlying fundamentals of Rydberg atoms, followed by the principles and schemes of RAQRs. Then, we overview the state-of-the-art studies from both physics and communication societies. Furthermore, we conceive Rydberg atomic quantum single-input single-output (RAQ-SISO) and multiple-input multiple-output (RAQ-MIMO) schemes for facilitating the integration of RAQRs with classical wireless systems. Finally, we conclude with a set of potent research directions.
@@ I Introduction
Next-generation wireless systems are expected to support a variety of functionalities and applications. Integrated communication and sensing functionalities of next-generation wireless require particularly high sensitivity and data rates, while supporting large-scale connectivity and ultra-low latency. Therefore, the radio frequency (RF) receivers should possess broadband tunability and wideband processing capability. Based upon advanced antennas and integrated circuit (IC), cutting-edge RF receivers support signal receptions by well-calibrated antennas, filters, amplifiers, and mixers [1]. The sensitivity is limited by the receiver’s bandwidth, noise figure, and the minimum signal-to-noise ratio (SNR) required to demodulate the desired signal. Given a specific sensitivity, strong RF reception capability can be attained by using multi-antenna based multiple-input multiple-output (MIMO) schemes and sophisticated signal processing methods to combat channel fading and noise contamination. To receive wideband RF signals at different frequencies, both broadband tunability and wideband processing capability may be realized by stacking several IC branches along with extra metallic components, where each branch is responsible for a specific frequency band. Such RF systems are however band-limited and bulky, power-thirsty, and suffer from numerous design challenges, especially for large-scale, high-frequency receivers.
@@
The Rydberg atomic quantum receiver (RAQR) concept emerges as a radical solution to these RF reception challenges [2, 3, 4, 5]. It relies on the creation of Rydberg states (excitation), which have remarkably high dipole moment and hence are eminently suitable for sensing the impinging RF signals. The response to an RF signal in RAQR is detected optically via change in the atomic spectroscopy signal. The RF field’s amplitude measurements are directly linked to the International System of Units (SI) and they set absolute (atomic) standards for the RF electric fields from a metrology standpoint. Additionally, they are capable of directly down-converting RF signals, spanning from direct-current (DC) to Terahertz (THz), to the baseband without using any mixer, hence significantly simplifying the receiver’s structure and yielding a compact form-factor relative to conventional counterparts.
@@
To fully unlock the potential of RAQRs in communication and sensing, we offer a panoramic overview of RAQRs. We commence by highlighting the fundamentals of Rydberg atoms and proceed to powerful receivers, including their principles, schemes, impairments, and their critical comparisons to antenna-based receivers. We then summarize their capabilities and distinctive features, surveying typical studies from diverse research communities. Furthermore, we conceive the Rydberg atomic quantum single-input single-output (RAQ-SISO) and RAQ-MIMO schemes to facilitate the integration of RAQRs with existing classical wireless systems. We finally forecast several promising future directions in this very first article on RAQRs in wireless communication and sensing from an engineering perspective.
@@
Figure 1: (a) A simplified diagram of atomic structure for ground-state and Rydberg atoms. (b) Scaling of the relevant properties of Rydberg states with the alkali atom principal number ($n$). (c) Radial probability distribution of Rydberg states of Cs atoms for $n=5,90$ ($l=2$). (d) The electron probability distribution $|\psi_{nlm}|^{2}$ for an atom excited to a state $n=5$, $l=2$, $m=1$ and for a Rydberg atom in state $n=90$, $l=2$, $m=1$, respectively.
@@ II Rydberg Atomic Quantum Receiver
RAQRs exploit the properties of Rydberg atoms, hence we commence by their brief overview, followed by their principles and schemes. Finally, we present the impairments of RAQRs, and compare them to classical RF receivers.
@@ II-A1 Atomic Structure [6]
Quantum mechanics provides a comprehensive description of the structure, properties and behavior of atoms. Each atom has a nucleus at its core and there are electrons that orbit around the nucleus. The energy of electrons in these orbits is quantized, i.e., they occupy only certain discrete energy-levels and cannot be traced in between these levels. However, electrons can transition from one energy level to another by absorbing or emitting a photon with energy equal to the difference between the energy levels. Electrons have a wave-particle dual nature. In the quantum mechanical description, the state of an electron is described by a wavefunction $\psi$, which contains all information about the electron’s position, energy, spin, orientation, etc. The square of the wavefunction, $|\psi|^{2}$ gives the probability density (i.e. probability per unit volume) of finding the electron in a certain spatial region around the nucleus. These regions are termed as orbitals. The unique quantum state of an electron is described by four quantum numbers:
@@
  • Azimuthal or orbital angular momentum quantum number $\bm{l}$ (non-negative integer $l\leq n-1$): It accounts for the sub-shell (sub-level) that the electron occupies, with each sub-shell having unique characteristic shapes.
@@
If mediated by photons of electromagnetic (EM) field, the transition of an electron from one quantum state to another is called an electric dipole transition. Based on the conservation of angular momentum and parity [^1] Parity refers to the symmetry of a wavefunction under spatial inversion., only those (electric dipole) transitions are allowed which satisfy the selection rules: ① $\Delta l=\pm 1$, ② $\Delta m=0,\pm 1$, and ③ the initial and final states must have opposite parities.
@@ II-A2 Properties of Rydberg State
Alkali atoms have single valence electrons in their ground state in the outermost orbital. This represents a relatively loose bound and therefore it can be excited to a high energy level associated with a high principal quantum number, of say $n>10$ for atomic systems. When such an excitation is created, the atom and its corresponding state are Rydberg ones. The atomic structures of the ground-state and Rydberg-state atoms are schematically shown in Fig. 1(a). The Rydberg states of any alkali atom and a hydrogen atom have essentially similar structures and properties. In both cases, a single electron occupies a high-$n$ state and has an effective core of unit positive charge at the center. The only difference lies in the finite size of the positively charged core, which is larger for alkali atoms relative to hydrogen atoms. This leads to a small correction (via quantum the defect parameter $\delta_{l}$) to the energies for non-hydrogenic atoms in particular for low $l$ states. But again, one can say that the general properties of all Rydberg atoms are quite similar. The scaling of the relevant properties of Rydberg states with $n$ is shown in Fig. 1(b). The $n$-dependence arises from the Rydberg atom wavefunction characteristics. For RAQRs, the typical choice of alkali atoms falls on either Caesium (Cs) or Rubidium (Rb) due to the availability of commercial laser systems for the transition wavelengths of these species.
[…]
@@ II-A3 Rydberg Atoms in Electric Fields
The sensitivity of Rydberg atoms to electric fields arises due to the high polarizability of Rydberg states. In the presence of electric field, the energy levels shift for different states in a way similar to the energy-level splitting of a hydrogen atom. This is called the Stark effect and the associated Stark shifts can be calculated by direct diagonalization of the Hamiltonian matrix. The plot of atomic energy levels as a function of the electric field strength applied is termed as the Stark map and such a map is shown in Fig. 2 for Cs atoms in the vicinity of the $30D$ state. The red dashed line shows the energy level in the absence of electric field. The low-$l$ ($l=0,1,2$) Rydberg states are non-degenerate and experience a second-order Stark shift, i.e., a quadratic scaling with electric field amplitude $\mathcal{E}$ given by $\Delta E_{\text{DC}}=\alpha\mathcal{E}^{2}/2$, where $\alpha$ is the DC polarizability of the Rydberg state. For $l>3$, the states are degenerate and therefore experience a linear Stark shift similar to the hydrogen Stark spectrum.
When the electric field interacting with atoms is constant or fluctuates at a low frequency, it is called DC Stark shift, as mentioned above. When a high-frequency electric field interacts with the atoms, it results in AC Stark shift, which has quadratic scaling with the field amplitude $\mathcal{E}$. The magnitude and sign of the AC Stark shift depends on the detuning $\Delta$ of the electric field from the resonant transition frequency and scales as $\Delta E_{\text{AC}}\propto 1/\Delta$. This implies that the magnitude of the shift is larger when the electric field is closer to resonance, while its sign is positive (negative) for blue-detuned (red-detuned) electric fields. The Stark effect causes coupling between closely-spaced energy levels leading to the mixing of energy eigenstates. As a result of mixing the of wavefunctions of two states anti-crossings appear in the Stark map, as a gap between two energy curves where they would otherwise cross each other. This has been explained in detail in [6].
II-B Principles and Schemes of RAQRs
@@ II-B1 Rydberg-Electromagnetically Induced Transparency (EIT) Spectroscopy and the Standard Scheme[5]
EIT is a quantum interference phenomenon, where two excitation pathways of a three-level system interfere to open a transmission window in the absorption spectrum of atomic resonance. This is typically fulfilled by counter-propagating a pair of laser beams, each of which corresponds to an excitation pathway. Since this is a coherent process, it is ultra-sensitive to phase perturbations and energy level shifts of the three-level system.
@@
Rydberg excitation is prepared via a two-photon excitation process, where a pair of counter-propagating laser beams of different wavelengths, namely the ‘probe’ and ‘coupling (control)’, are overlapped inside an alkali vapor cell, as shown in Fig. 3(a). This is accomplished via Rydberg-EIT in a ladder-type configuration, namely |$\ket{1}$$\rightarrow$|$\ket{2}$$\rightarrow$|$\ket{3}$$\rightarrow$|$\ket{4}$ (‘standard’) in Fig. 3(h). The probe beam excites the atom from its ground state to an intermediate excited state, while the coupling beam takes it to a high-energy (principal number $n$) state, namely the Rydberg state. When an external RF signal’s frequency is (near) resonant with two adjacent Rydberg states, changes can be observed in the optical spectroscopic signal in the following three scenarios, namely (i) probe beam only, (ii) probe & coupling beams, and (iii) probe & coupling beams plus RF signal, as illustrated in Fig. 3(b) using realistic experimental parameters. With a further increase in the RF field strength, the Rydberg-EIT peak splits into two, known as the Autler-Townes splitting (ATS). The RF signal strength can be determined by measuring the frequency splitting of the two peaks of ATS.
@@
This is the standard scheme (structure) used in RAQRs. The electron transitions take place among the four energy levels (quantum states) |$\ket{1}$, |$\ket{2}$, |$\ket{3}$, and |$\ket{4}$ with resonant transition frequencies $\omega_{(P,C,s)}$ and detunings $\Delta_{(P,C)}$, as shown in Fig. 3(h). Each quantum state has its own decay rate, denoted by $\gamma$. The transition |$\ket{1}\rightarrow\ket{2}$ is typically chosen to be D2 transition of Cs (Rb). |$\ket{3}$ could be any Rydberg state, and the excitation is created by tuning the coupling laser wavelength to the appropriate value. Finally, |$\ket{4}$ in the ‘standard’ box of Fig. 3(h) is the neighboring Rydberg state and its choice is governed by the particular RF frequency that the RAQR would be set up to receive. In this scheme, the measurement sensitivity approaches $\sim$ µV/cm/$\sqrt{\text{Hz}}$ [2].
II-B2 Superheterodyne Scheme for Enhanced Sensitivity and Phase Extraction [7, 8]
RAQR’s measurement sensitivity can be further improved by exploiting the superheterodyne principle, where a local oscillator (LO) as a reference is added to the structure for detecting relatively weak RF signals. This scheme is presented in Fig. 3(c). The RF signal to be detected can be at intermediate frequency within the instantaneous bandwidth of RAQRs. The LO facilitates the creating of microwave-dressed Rydberg states and the frequency splitting measurement applied in the standard scheme is converted into an amplitude-modulated measurement in the presence of these two RF signals. As such, Rydberg atoms serve as the role of a conventional RF mixer. In this scheme, the phase of the desired RF signal can be measured relative to that of the LO from the modulated amplitude of the probe beam. The sensitivity has been improved to 55 nV/cm/$\sqrt{\text{Hz}}$ [8].
II-B3 Optical Laser Array (OLA) for Enhanced Sensitivity [9]
The sensitivity of the superheterodyne scheme can be further enhanced by exploiting the OLA structure (Fig. 3(d)), where multiple laser beams are generated in parallel by employing well-designed gratings (split-beam and collimator gratings). Therefore, multiple parallel Rydberg atomic regions are created for simultaneously receiving RF signals. The multiple laser beams are combined and detected by the photodetector (PD) for realizing high-sensitivity detectors. More particularly, the sensitivity has been enhanced to 19 nV/cm/$\sqrt{\text{Hz}}$ [9]. By employing the OLA scheme to the most advanced Rydberg sensor, the sensitivity of $<$ 1 nV/cm/$\sqrt{\text{Hz}}$ is achievable [9].
II-B4 Polarization-Sensitive Scheme (PSS) for Vector Field Sensing [10]
The polarization of RF signals can be determined by RAQRs by exploiting three independent and orthogonally-polarized LOs. Specifically, these LOs are $x$-, $y$-, and $z$-polarized, respectively, as seen in Fig. 3(e). The ladder-type scheme is presented in the ‘PSS’ box of Fig. 3(h). The RF signal to be detected is mixed with these LO signals and thus three signal beats are created for extracting their relative phases. Notably, these relative phases facilitate the determination of the signal polarization.
Figure 3: (a) The standard scheme, and (b) Rydberg-EIT spectroscopy signal and ATS. (c) The superheterodyne scheme, (d) OLA-based scheme, (e) PSS scheme, (f) continuous-band detection scheme, and (g) multiband detection schemes. (h) The ladder-type scheme for showing the electron transition of the above schemes.
II-B5 Adjacent Rydberg Resonance Tuning (ARRT) for Enhanced Instantaneous Bandwidth and Continuous-Band Detection [11, 12]
The RF signal to be detected is typically limited to a narrow instantaneous bandwidth ($\leq$ 10 MHz [5]) around the transition frequency of coupled Rydberg states. The reason for this is that a wide bandwidth leads to substantial RF detuning, creating asymmetric ATS peaks. Once the RF detuning exceeds some narrow range, the RF signal is no longer resonant with the Rydberg state and the ATS cannot be observed. Extending the instantaneous bandwidth is feasible by exploiting the ARRT scheme of [11], where an auxiliary RF signal is employed to couple another Rydberg state adjacent to that coupled with the desired RF signal, as shown in Fig. 3(f) and the ‘ARRT’ box of Fig. 3(h). The instantaneous bandwidth can be up to $100$ MHz [3] and reference [39] of [3]. This scheme facilitates the detection of RF signals distributed across a continuous frequency range (e.g., $\geq 1$ GHz [12]).
II-B6 Cascade Rydberg States (CRS) [13], Different Rydberg Final States (DRFS) [14], and Mixed Atomic Species (MAS) [15] for Simultaneous Multiband Detection
RAQRs are capable of simultaneously detecting multiple RF signals having different carrier frequencies by exploiting three different schemes, as shown in Fig. 3(g) and the ‘CRS’, ‘DRFS’, ‘MAS’ boxes of Fig. 3(h). Specifically, the CRS scheme [13] exploits a series of high-orbital-angular-momentum states that are simultaneously cascaded by multiple RF signals having their carrier frequencies in descending order (from THz to MHz). Furthermore, the DRFS scheme [14] simultaneously harnesses multiple transitions from an identical Rydberg state to multiple different Rydberg states, where each transition corresponds to the RF signal having a specific frequency. Lastly, the MAS scheme [15] realizes the simultaneous detection by exploiting mixed alkali atomic species (Cs and Rb). Each atom species is excited relying on a pair of laser beams having specific wavelengths and hence detecting the corresponding RF signal having a specific frequency.
II-B7 Other Promising Schemes
Beyond the above schemes, a suite of more sophisticated schemes emerge. For instance, by exploiting the machine learning models [16], the RAQRs can be trained to better receive RF signals than those schemes relying on the master equation. Additionally, the scheme exploiting the many-body effect exhibits an enhanced sensitivity, as seen in [3] and reference [124] therein, paving the way for approaching the standard quantum limit (SQL). Furthermore, other advanced schemes exploiting Schrodinger cat states and squeezed states may potentially allow RAQRs to surpass the SQL, as discussed in [4] and references [101], [103] therein.